OpenStudy (anonymous):

5/8 divided by 15/24 A. 1 B. 3 C. 5/3 D. 3/5

OpenStudy (anonymous):

Nevermind... Its a

OpenStudy (paki):

(5/8) / (15/24) = 5*24 / 8*15 = solve now...